how much water i have to add so that my 70% concentrated H.Cl solution becomes 50% concentrated solution ?


Answer by stanbon(75887)

how much water i have to add so that my 70% concentrated H.Cl solution becomes 50% concentrated solution ?
------
Equation::
Assuming you have 20ml of 70% HCL,
water + water = water
0.30*20 + 1*x = 0.50(20+x)
-----
30*20 + 100x = 50*20 + 50x
50x = 20*20
50x = 400
x = 8 ml
=====================
Cheers,
Stan H.
===================